will you please solve this problem for me it will be greatly appreciated:
solve: (n)/(n-3) + n=(7n-18)/(n-3)

Answer by stanbon(48532)

solve: (n)/(n-3) + n=(7n-18)/(n-3)
Multiply thru by (n-3) to get:
n + n(n-3) = 7n-18
n + n^2 - 3n = 7n - 18
n^2 -9n +18 = 0
(n-6)(n-3) = 0
n = 6
--------------
n cannot be 3 because that would make the
denominator of the original prolem be zero.
===============
Cheers,
Stan H.